OpenStudy (anonymous):

5.In the following equation, identify the x-intercepts in the graph y = 3x^2 + 19x – 40. I'm stuck on this problem can anyone help? :)

mathslover (mathslover):

You mean to say that we have to find x intercept?

OpenStudy (anonymous):

Where are you having problem??

mathslover (mathslover):

Ok so just put y =0 : \[\large{0=3x^2+19x-40}\] \[\large{0=3x^2+24x-5x-40}\] \[\large{0=3x(x+8)-5(x+8)}\] \[\large{0=(3x-5)(x+8)}\]

OpenStudy (anonymous):

For finding x intercept we put y = 0..

OpenStudy (anonymous):

Oh okay, sry I forgot about that and yeah had trouble factoring it and now I see. This means our x= -8 and 5/3

OpenStudy (anonymous):

Yep..

mathslover (mathslover):

right

mathslover (mathslover):

3x-5=0 x+8=0 3x=5 x=-8 x=5/3

OpenStudy (anonymous):

Right, I'd also like to add that in case you forget whether to make the y or the x zero when it comes to either intercept remember that any x-intercept would have to be (x, 0) and that any y-intercept would have to be (0, y). Cheers

OpenStudy (anonymous):

K thx a ton :)